OpenStudy (sugarmochi):

asdfghjkl h e l p Solve: 32x^2 - 40x - 12 = -8x + 12

OpenStudy (idkanything):

x=-0.5

OpenStudy (sugarmochi):

sorry- but i forgot to mention its quadratic equations :< and there's supposed to be two answers-- i t h i n k

OpenStudy (idkanything):

oh then idk @KJSaif

KjSaif (kjsaif):

\[x=\frac{ -1 }{ 2 }\\] and\[x=\frac{ 3 }{ 2 }\]

KjSaif (kjsaif):

-1/2 and 3/2

OpenStudy (sugarmochi):

thankyou ! But how do i exactly reach the the answers...?

KjSaif (kjsaif):

im glad you asked :)

KjSaif (kjsaif):

Step 1: Subtract -8x+12 from both sides. 32x2−40x−12−(−8x+12)=−8x+12−(−8x+12) 32x2−32x−24=0 Step 2: Factor left side of equation. 8(2x+1)(2x−3)=0 Step 3: Set factors equal to 0. 2x+1=0 or 2x−3=0

OpenStudy (idkanything):

good job kjsaif!

KjSaif (kjsaif):

Thanks @Idkanything

OpenStudy (idkanything):

i just knew only 1 answer

OpenStudy (idkanything):

@KJSaif